# Physics Informed Generative Machine Learning for Accelerated Quantum-centric Supercomputing

### v1

@@ VI Conclusion and Future Outlook
In this paper, we propose PIGen-SQD, a novel QCSC workflow based on SQD framework that leverages a strategic orchestration of sampling efficiency of quantum computers, many-body intuitions and generative power of machine learning to provide better convergence with substantially reduced classical resource utilization. QCSC/ SQD utilizes the quantum computer as a sampling engine to generate configuration subspace in which the molecular Hamiltonian is projected and diagonalized. In the presence of quantum hardware noise, SQD requires a configuration recovery module, a heuristic technique that “corrects” the noise-affected samples with iterative filtering based upon the correct spin-particle number symmetry and average occupation probability. Accuracy of SQD relies upon the the degree of initial overlap with the target state provided by the sampled distribution and the capability of heuristic configuration recovery scheme to predict dominant configurations. However, the key challenge is that the state preparation on current quantum hardware does not necessarily guarantee high quality samples to provide this support to the target state and may require prohibitively huge number of shots which collectively adds a heuristic component to the SQD pipeline. To address this, here we introduce a physics-informed pre-processing which employs perturbative measures with implicit low rank tensor decompositions to identify a particular group of dominant configurations. These perturbative measure when combined with hardware samples, yields a substantially improved initial overlap to the target state providing a head-start for configuration recovery. This is followed by integration of generative ML models (RBM in this case) which upon training with this sample distribution, generates new configurations while exploring only the relevant sector of the Hilbert space of chemical relevance. Our numerical demonstrations with 24, 32 and 40 qubit molecular simulations conducted on state-of-the-art IBM Heron R2 processors suggest PIGen-SQD outperforms traditional SQD in both accuracy and efficiency, providing up to one-order of better energy accuracy with up to 70 reduction in the corresponding subspace dimension required for Hamiltonian diagonalization. Such accuracy with substantially reduced HPC resource requirements is observed even under strong electronic correlations where SQD struggles to predict chemically accurate results.
@@
From a near-future application perspective PIGen-SQD can be integrated with other SQD variants such as Sample-based Krylov Quantum Diagonalization (SKQD)[38] or its more efficient variant using quantum stochastic drift protocol (SqDRIFT)[39] for guaranteed and faster convergence along with enhanced computational efficiency. Moreover, the current architecture have huge scope of technical improvements such as RBM can be replaced with more advanced neural network models such as transformers which has recently been shown to provide unique advantages while dealing with complex molecular systems. With continued improvements in quantum-hardware capabilities and machine-learning components, PIGen-SQD presents a promising pathway toward scalable molecular simulations, advancing the field toward resolving several longstanding challenges in chemistry.
Acknowledgments
AUTHOR DECLARATIONS
Conflict of Interest:

### v2

@@ VI Conclusion and Future Outlook
In this paper, we propose PIGen-SQD, a novel QCSC workflow based on SQD framework that leverages a strategic orchestration of sampling efficiency of quantum computers, many-body intuitions and generative power of machine learning to provide better convergence with substantially reduced classical resource utilization. QCSC/ SQD utilizes the quantum computer as a sampling engine to generate configuration subspace in which the molecular Hamiltonian is projected and diagonalized. In the presence of quantum hardware noise, SQD requires a configuration recovery module, a heuristic technique that “corrects” the noise-affected samples with iterative filtering based upon the correct spin-particle number symmetry and average occupation probability. Accuracy of SQD relies upon the the degree of initial overlap with the target state provided by the sampled distribution and the capability of heuristic configuration recovery scheme to predict dominant configurations. However, the key challenge is that the state preparation on current quantum hardware does not necessarily guarantee high quality samples to provide this support to the target state and may require prohibitively huge number of shots which collectively adds a heuristic component to the SQD pipeline. To address this, here we introduce a physics-informed pre-processing which employs perturbative measures with implicit low rank tensor decompositions to identify a particular group of dominant configurations. These perturbative measure when combined with hardware samples, yields a substantially improved initial overlap to the target state providing a head-start for configuration recovery. This is followed by integration of generative ML models (RBM in this case) which upon training with this sample distribution, generates new configurations while exploring only the relevant sector of the Hilbert space of chemical relevance. Our numerical demonstrations with 24, 32 and 40 qubit molecular simulations conducted on state-of-the-art IBM Heron R2 processors suggest PIGen-SQD outperforms traditional SQD in both accuracy and efficiency, providing up to one-order of better energy accuracy with up to 90 reduction in the corresponding subspace dimension required for Hamiltonian diagonalization. Such accuracy with substantially reduced HPC resource requirements is observed even under strong electronic correlations where SQD struggles to predict chemically accurate results.
@@
From a near-future application perspective PIGen-SQD can be integrated with other SQD variants such as Sample-based Krylov Quantum Diagonalization (SKQD)[44] or its more efficient variant using quantum stochastic drift protocol (SqDRIFT)[47] for guaranteed and faster convergence along with enhanced computational efficiency. Moreover, the current architecture have huge scope of technical improvements such as RBM can be replaced with more advanced neural network models such as transformers[72] which has recently been shown to provide unique advantages while dealing with complex molecular systems. With continued improvements in quantum-hardware capabilities and machine-learning components, PIGen-SQD presents a promising pathway toward scalable molecular simulations, advancing the field toward resolving several longstanding challenges in chemistry.
Author contributions
Conceptualization: Chayan Patra, Rahul Maitra
Methodology: Chayan Patra, Rahul Maitra
Investigation: Chayan Patra, Rahul Maitra, Dibyendu Mondal
Formal analysis: Chayan Patra, Rahul Maitra, Dibyendu Mondal
Data curation: Chayan Patra
Writing - original draft: Chayan Patra, Rahul Maitra
Writing - review and editing: All authors
Software: Chayan Patra, Mostafizur Rahaman Laskar, Sonaldeep Halder, Richa Goel
Visualization: Chayan Patra
Supervision: Rahul Maitra
Funding acquisition: Rahul Maitra
Conflicts of interest